OpenStudy (anonymous):

x²-3x-10=0 need really fast help

OpenStudy (ajprincess):

Pls tell me two factors of -10 so that their sum is -3.

OpenStudy (anonymous):

5 and 2 one negatve one positive

OpenStudy (anonymous):

-5 and 2

OpenStudy (ajprincess):

yup good:) Nw -3x can be written as -5x+2x. Agree?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

good:) \(x^2-3x-10=0\) \(x^2-5x+2x-10=0\) what do u get when u factorise \(x^2-5x\)?

OpenStudy (anonymous):

ughhhhhh im stuck lol

OpenStudy (ajprincess):

\(x^2-5x=x*x-5*x\) What is common to both \(x^2\) and \(5x\)?

OpenStudy (anonymous):

x

OpenStudy (ajprincess):

good:) nw when u take it out what remains

OpenStudy (anonymous):

x-5

OpenStudy (anonymous):

aj?

OpenStudy (ajprincess):

ya. right:) So \(x^2-5x=x(x-5)\) Nw can u factorise 2x-10

OpenStudy (anonymous):

x-5

OpenStudy (ajprincess):

well 2x-10=2(x-5) Nw can u factorise x(x-5)+2(x-5)

OpenStudy (ajprincess):

@kittenbalm

OpenStudy (anonymous):

ugh how do i foil?

OpenStudy (ajprincess):

treat (x-5) as one term. for simplicity let us assume (x-5)=y. xy+2y. can u factorise this?

OpenStudy (anonymous):

yes you can

OpenStudy (ajprincess):

I was asking whether u can? I can

OpenStudy (anonymous):

oh no

OpenStudy (ajprincess):

xy+2y What is common to both xy and 2y?

OpenStudy (anonymous):

y

OpenStudy (ajprincess):

good:) So what remains when u take it out?

OpenStudy (anonymous):

x+2

OpenStudy (ajprincess):

right:) xy+2y=y(x+2) our y=x-5. so x(x-5)+2(x-5)=(x-5)(x+2) \(x^2-3x-10=0\) \(x^2-5x+2x-10=0\) \(x(x-5)+2(x-5)=0\) \((x-5)(x+2)=0\) x-5=0 or x+2=0 Hope u can find the two values of x nw:)

OpenStudy (anonymous):

-5 and 2

OpenStudy (ajprincess):

Nope. pls check again:)

OpenStudy (anonymous):

oh 5 and -2

OpenStudy (ajprincess):

yup good:)